the length of a rectangle is 4 cm greater than its width. If the length and the width were increased by 3 cm, the area of the rectangle would be in creased by 93 cm. What are the dimensions fo the original rectangle?
Answer by lwsshak3(11628)

he length of a rectangle is 4 cm greater than its width. If the length and the width were increased by 3 cm, the area of the rectangle would be in creased by 93 cm. What are the dimensions fo the original rectangle?
**
let w=width of original rectangle
w+4=length of original rectangle
..
w+3=width of enlarged rectangle
w+4+3=w+7=length of enlarged rectangle
..
area=length*width
(w+3)(w+7)-w(w+4)=93
w^2+10w+21-w^2-4w=93
6w=72
w=12
w+4=16
ans:
width of original rectangle=12 cm
length of original rectangle=16 cm
..